# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ E.1 Algorithm Pseudocode
2:Data-Preprocessing: Compute local neighborhood means using Nearest Neighbor Algorithm and Ligand-Receptor features $f_{\text{LR}}$ using LIANA+ $\triangleright$ As defined in Equation 9 and 10
@@
E.2 Time Complexity Analysis
The training time of ContextFlow is comparable to that of Minibatch-OT FM (Tong et al., 2024) (Figure 3), as both methods solve an entropic variant of optimal transport using the GPU-optimized Sinkhorn algorithm, alongside forward and backward propagation steps that are also GPU-accelerated. Although ContextFlow incorporates prior knowledge, such as spatial smoothness (Equation 9) and cell-cell communication patterns (Equation 10), their corresponding features are computed once during preprocessing, resulting in a one-time cost. The precomputed features can be reused across multiple hyperparameter settings and model variants, making ContextFlow highly scalable and efficient.
Data Preprocessing. The following preprocessing steps generate additional biologically informed features that complement the original transcriptomic profiles. These features incur a one-time computational cost and can be reused across different experiments and model configurations.
Spatial Smoothness (SS). We employ a nearest neighbor (NN) algorithm for calculating the mean of local transcriptomic features for each cell. The computational complexity of the NN search is known to be $O(N^{2}d)$, where $N$ denotes the total number of points considered and $d$ denotes the data dimension.
Cell-Cell Communication Patterns (LR). We employ spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), for computing LR features, where we applied the cosine similarity metric to gene expression profiles and used the recommended hyperparameters. The exact runtime complexity for LIANA+ is unknown. Table 6 summarizes the total time taken for the Brain Regeneration and Mouse Organogenesis datasets in our case.
@@ E.2 Time Complexity Analysis
Brain Regeneration (Wei et al., 2022) 28,780 23.35
Mouse Organogenesis (Chen et al., 2022) 399,248 200.40
Training of ContextFlow. The training time complexity largely depends on the total number of training epochs ($E$), mini-batch size ($B$), time per forward and backward pass ($P$), transcriptomic feature dimension ($d$), and total LR pairs ($l$). Below, we compute the time complexity for each step in ContextFlow:
TPM Construction. The construction of the transition plausibility matrix involves a calculation of pairwise distances for each mini-batch, resulting in the runtime of $O(B^{2}(l+d))$.
Sinkhorn Algorithm. According to Theorem 3.1, we know that Sinkhorn iterations can be adapted to solve the prior-aware entropy regularization problem (Equation 13). Since the Sinkhorn algorithm has a well-known quadratic time complexity (Cuturi, 2013), the runtime for computing minibatch OT couplings in ContextFlow is $O(B^{2})$.
Figure 3: Training time comparisons for different algorithms with a mini-batch size of $256$ on the Brain Regeneration dataset.
Total Runtime. Putting pieces together, across all the training epochs, the total runtime complexity of ContextFlow turns out to be $O(E\times(B^{2}(d+l)+P))$. As shown in Figure 3, the runtime is linearly dependent on the total epochs $E$, with different linear rates for various configurations. CFM is the fastest because it bypasses the optimal transport coupling step required by the other methods.
@@ Appendix F Spatiotemporal Optimal Transport
In this section, we compare the recent state-of-the-art spatiotemporal alignment methods, including DeST-OT (Halmos et al., 2025) and TOAST (Ceccarelli et al., 2025), with our prior-aware entropy regularized (PAER) OT objective used in ContextFlow (CTF-H). It is important to note that these OT methods are not generative models and are only used for pairwise alignment tasks. ContextFlow, on the other hand, is a generative model that learns a dynamic flow across the time horizon and utilizes OT couplings to design better conditional paths for regression. In particular, we compute metrics described in DeST-OT on the Axolotl Brain Regeneration dataset following the same setup as used in flow matching. Specifically, for each time step, we randomly sample a batch of $1000$ cells and compute the corresponding coupling matrix $\mathbf{\Pi}$, which is then used to derive the metrics. We use the CTF-H ($\lambda=0.8$) version of ContextFlow for comparison.
F.1 Metric Comparison
@@
DeST-OT introduces an OT objective for aligning spatial transcriptomic tissue slices from different developmental timesteps, with an emphasis on modeling cell growth and tissue expansion/contraction. The growth distortion metric is designed to assess whether the inferred growth pattern aligns with the changes in cell-type abundance across timesteps. As shown in Table 7, for the growth distortion metric, we find that our CTF-H OT is competitive with DeST-OT and TOAST, despite DeST-OT being specifically developed with consideration for cell growth.
Table 7: Comparison on growth distortion.
Static Pair DeST-OT TOAST CTF-H OT
$1/2$ $0.0007$ $0.0000$ $0.0000$
$2/3$ $0.0042$ $0.0000$ $0.0000$
$3/4$ $0.0027$ $0.0000$ $0.0000$
$4/5$ $0.0009$ $0.0000$ $0.0000$
@@
The migration metric is another important metric introduced in DeST-OT, which measures whether the coupling results in realistic cell movements between timesteps. As seen in Table 8, DeST-OT achieves the best performance, highlighting the advantage of its growth-aware objective compared to TOAST and CTF-H OT, which do not explicitly model tissue expansion or contraction.
Table 8: Comparison on migration.
Static Pair DeST-OT TOAST CTF-H OT
$1/2$ $308.97$ $793.71$ $780.16$
$2/3$ $551.53$ $1103.29$ $1037.29$
$3/4$ $1015.73$ $2052.65$ $1991.43$
$4/5$ $777.13$ $2257.02$ $2222.07$
@@
Lastly, we compute how similar the transcriptomic values of coupled cells are using a Coupled Transcriptomic Distance metric, which is defined as $\sum_{k=1}^{N}\sum_{l=1}^{M}\big\|X_{t_{i}}[k,:]-X_{t_{i+1}}[l,:]\big\|^{2}\times\Pi_{i,j}$, where $X_{t_{i}}[k,:]$ represents the transcriptomic feature of cell $k$ from timestep $t_{i}$ and $X_{t_{i+1}}[l,:]$ represents the transcriptomic feature of cell $l$ from timestep $t_{i+1}$, and $\Pi$ is the OT coupling matrix. From Table 9, we can observe that CTF-H OT is competitive with both DeST and TOAST.
Table 9: On coupled transcriptomic distance.
Static Pair DeST-OT TOAST CTF-H OT
$1/2$ $32.64$ $34.13$ $33.58$
$2/3$ $14.22$ $21.87$ $18.01$
$3/4$ $42.26$ $43.89$ $43.26$
$4/5$ $18.47$ $20.54$ $20.14$
F.2 Runtime Analysis with Varying Sample Size
We also compare the runtime complexity of the above-mentioned OT methods, as shown in Table 10. CTF-H OT is the fastest of the three, followed by DeST-OT and TOAST, and is competitive in the metrics above. We also observe that DeST-OT is the slowest, as expected, since its OT objective includes a Gromov-Wasserstein term, which has $O(n^{3})$ runtime, along with other growth- and tissue-distortion-specific terms.
@@ F.2 Runtime Analysis with Varying Sample Size
Table 10: Runtime (s) with varying sample size.
Sample Size DeST-OT TOAST CTF-OT
$10$ $0.0111$ $0.0227$ $0.0009$
$50$ $0.0342$ $0.0114$ $0.0018$
$100$ $0.1892$ $0.0167$ $0.0024$
$150$ $0.4035$ $0.0247$ $0.0053$
$200$ $0.4913$ $0.0337$ $0.0059$
$250$ $0.4571$ $0.0426$ $0.0061$
$300$ $0.6252$ $0.0543$ $0.0074$
$350$ $0.6974$ $0.0656$ $0.0078$
$400$ $0.8612$ $0.0817$ $0.0117$
$450$ $1.1028$ $0.0983$ $0.0110$
$500$ $1.5478$ $0.1197$ $0.0107$
$550$ $1.7448$ $0.1468$ $0.0165$
$600$ $1.9295$ $0.1905$ $0.0209$
$650$ $2.1282$ $0.2077$ $0.0201$
$700$ $2.7013$ $0.2309$ $0.0235$
$750$ $3.2951$ $0.2574$ $0.0327$
$800$ $4.0964$ $0.3001$ $0.0382$
$850$ $4.2001$ $0.3229$ $0.0339$
$900$ $4.4582$ $0.3798$ $0.0483$
$950$ $5.0965$ $0.4206$ $0.0509$
$1000$ $6.1452$ $0.4931$ $0.0375$
Tables 7-10 demonstrate that the design choices of ContextFlow enable it to be highly scalable compared to state-of-the-art spatiotemporal alignment methods, while remaining competitive across several spatiotemporal OT alignment metrics.
(a) MOTFM Implausible Transitions
(b) ContextFlow Implausible Transitions
Figure 4: Comparison of biologically implausible cell type couplings between Stage 0 and Stage 1 of the Brain Regeneration Dataset (Wei et al., 2022), under the Entropic-OT and ContextFlow Regularized-OT formulations. Biological implausibility is defined here as transitions involving excitatory–inhibitory lineage switches. Our formulation produces substantially fewer biologically implausible couplings (24) compared to MOTFM (54).
Appendix G Biological Priors and Dataset Visualizations
@@
To demonstrate the need of integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework. From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types. Figures 4(a) and 4(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings. Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
In our transport plan couplings, we observed the following cell type lineage switches: (i) Immature MSN $\rightarrow$ Immature nptxEX, (ii) Immature MSN $\rightarrow$ Immature dpEX, (iii) Immature MSN $\rightarrow$ Immature CMPN, (iv) Immature nptxEX $\rightarrow$ Immature cckIN, and (v) Immature nptxEX $\rightarrow$ Immature MSN
Of these, $54$ implausible transitions arose from the Entropic-OT plan, compared to $24$ from the PAER-OT plan, with the specific transitions detailed in the figure legends. We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right. In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information. These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
G.2 Cell type distributions over time
@@ G.2.3 Liver Regeneration
G.3 Ligand Receptor Interactions
Figure 8 shows the ligand-receptor score of the NPTX2-NPTXR pair in two consecutive slides from the Brain regeneration dataset (Wei et al., 2022). Similar activities are visible bilaterally in the cerebral cortex, suggesting that ligand–receptor interactions are preserved across time and spatially aligned with underlying tissue structure. This observation provides strong evidence that including LR interactions as contextual priors is biologically meaningful, as they capture functional communication signals between cells that remain stable across short time intervals.
@@ G.3 Ligand Receptor Interactions
Based on the activation of NPTX2–NPTXR in Figure 8, we observe that the corresponding communication pattern naturally biases the optimal couplings towards transitions such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 9). These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context. This example highlights the richness of the contextual information captured by our proposed biological prior and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process yields more interpretable, biologically consistent trajectories.

### v3

@@ E.1 Algorithm Pseudocode
2:Data-Preprocessing: Compute local neighborhood means using Nearest Neighbor Algorithm and Ligand-Receptor features $f_{\text{LR}}$ using LIANA+ $\triangleright$ As defined in Equation 4 and 5
@@
E.2 Training Details and Hyperparameters
Similar neural network architectures and training settings are used for all the evaluated baselines and ContextFlow. Specifically, we list the training hyperparameters used in our experiments:
  • Learning rate: $0.0005$
  • Learning rate scheduler: Cosine
  • Epochs: $10000$
  • NN: MLP
  • Depth: 44
  • Width: $100$
  • Radius of SS prior: $50$
  • LIANA+ [Dimitrov et al., 2024] settings for LR prior:
    • Resource: Consensus
    • nzprop: $0.01$
    • Similarity Metric: Cosine
  • Hyperparameter $\epsilon$ used in EOT formulation (Equation 2):
    • $0.1$ for Brain Regeneration experiments
    • $10$ for Organogenesis and Liver regeneration
E.3 Time Complexity Analysis
The training time of ContextFlow is comparable to that of Minibatch-OT FM [Tong et al., 2024] (Figure 4), as both methods solve an entropic variant of optimal transport using the GPU-optimized Sinkhorn algorithm, alongside forward and backward propagation steps that are also GPU-accelerated. Although ContextFlow incorporates prior knowledge, such as spatial smoothness (Equation 4) and cell-cell communication patterns (Equation 5), their corresponding features are computed once during preprocessing, resulting in a one-time cost. The precomputed features can be reused across multiple hyperparameter settings and model variants, making ContextFlow highly scalable and efficient.
Data Preprocessing. The preprocessing steps generate additional biologically informed features that complement the original transcriptomic profiles. These features incur a one-time computational cost and can be reused across different experiments and model configurations.
For the SS prior, we use a nearest-neighbor (NN) algorithm to compute the mean of local transcriptomic features for each cell. The computational complexity of the NN search is known to be $O(N^{2}d)$, where $N$ denotes the total number of points considered and $d$ denotes the data dimension. For the LR prior, we employ spatially informed bivariate statistics implemented in LIANA+ [Dimitrov et al., 2024], using the cosine similarity metric on gene expression profiles and the recommended hyperparameters. The exact runtime complexity for LIANA+ is unknown. Table 5 summarizes the total time taken for the Brain Regeneration and Mouse Organogenesis datasets.
@@ E.3 Time Complexity Analysis
Brain Regeneration [Wei et al., 2022] 28,78028{,}780 $23.35$
Mouse Organogenesis [Chen et al., 2022] 399,248399{,}248 $200.40$
Figure 4: Runtime comparisons across various algorithms on Brain Regeneration with batch size $256$.
Training of ContextFlow. The time complexity depends on the total number of training epochs ($E$), mini-batch size ($B$), time per forward and backward pass ($P$), transcriptomic feature dimension ($d$), and the total number of LR pairs ($l$). Below, we compute the time complexity for each step.
The construction of the transition plausibility matrix involves a calculation of pairwise distances for each mini-batch, resulting in the runtime of $O(B^{2}(l+d))$. According to Theorem 4.1, we know that Sinkhorn iterations can be adapted to solve the prior-aware entropy regularization problem (Equation 8). Since the Sinkhorn algorithm has a well-known quadratic time complexity [Cuturi, 2013], the runtime for computing minibatch OT couplings in ContextFlow is $O(B^{2})$. Putting pieces together, across all the training epochs, the total runtime complexity of ContextFlow turns out to be $O(E\times(B^{2}(d+l)+P))$. As shown in Figure 4, the runtime is linearly dependent on the total epochs $E$, with different linear rates for various configurations. CFM is the fastest because it bypasses the optimal transport coupling step required by the other methods.
@@ Appendix F Spatiotemporal Optimal Transport
In this section, we compare state-of-the-art spatiotemporal alignment methods, including DeST-OT [Halmos et al., 2025], MOSCOT [Klein et al., 2025] and TOAST [Ceccarelli et al., 2026], with our PAER-OT objective used in the entropic variant of ContextFlow (CTF-H). It is important to note that these OT methods are not generative models and are only used for pairwise alignment tasks. Our method, on the other hand, is a generative model that learns a dynamic flow across the time horizon and utilizes OT couplings to design better conditional paths for regression. In particular, we compute the metrics described in DeST-OT on the Axolotl Brain Regeneration dataset, using the same setup as in flow matching. Specifically, for each time step, we randomly sample a batch of $1000$ cells and compute the corresponding coupling matrix $\mathbf{\Pi}$, which is then used to derive the metrics. We use the CTF-H ($\lambda=0.8$) version of ContextFlow for comparison.
Table 6: Comparisons in terms of growth distortion metric.
Static Pair CTF-H OT DeST-OT TOAST MOSCOT
$(1,2)$ $0.0007$ $0.0000$ $0.0000$ $0.0000$
$(2,3)$ $0.0042$ $0.0000$ $0.0000$ $0.0000$
$(3,4)$ $0.0027$ $0.0000$ $0.0000$ $0.0000$
$(4,5)$ $0.0009$ $0.0000$ $0.0000$ $0.0000$
Table 7: Comparisons in terms of migration metric.
Static Pair CTF-H OT DeST-OT TOAST MOSCOT
$(1,2)$ $231.39$ $308.96$ $793.70$ $232.83$
$(2,3)$ $289.54$ $551.52$ $1103.28$ $395.15$
$(3,4)$ $2094.11$ $1015.72$ $2052.65$ $950.87$
$(4,5)$ $1763.51$ $777.13$ $2257.01$ $536.65$
@@
Growth Distortion. DeST-OT introduces an OT objective for aligning spatial transcriptomic tissue slices from different developmental timesteps, with an emphasis on modeling cell growth and tissue expansion/contraction. The growth distortion metric assesses whether the inferred growth pattern aligns with changes in cell-type abundance across timesteps. Table 6 shows that our CTF-H OT method is competitive, despite DeST-OT being specifically designed with cell growth in mind.
@@
Migration. The migration metric introduced in DeST-OT assesses whether the coupling produces realistic cell movements between timesteps. As seen in Table 7, DeST-OT and MOSCOT achieve the best performance overall, highlighting the advantage of its growth-aware objective compared to TOAST and CTF-H OT, which do not explicitly model tissue expansion or contraction. Nevertheless, CTF-H OT achieves superior performance, especially for the first two developmental stage pairs.
Coupled Transcriptomic Distance. Finally, coupled transcriptomic distance computes how similar the transcriptomic values of coupled cells are, which is defined as:
$\sum_{k=1}^{N}\sum_{l=1}^{M}\big\|X_{t_{i}}[k,:]-X_{t_{i+1}}[l,:]\big\|^{2}\times\Pi_{i,j},$
@@
where $X_{t_{i}}[k,:]$ represents the transcriptomic feature of cell $k$ from timestep $t_{i}$ and $X_{t_{i+1}}[l,:]$ represents the transcriptomic feature of cell $l$ from timestep $t_{i+1}$, and $\Pi$ is the OT coupling matrix. From Table 8, we can observe that CTF-H OT is competitive with DeST, MOSCOT, and TOAST.
Table 8: Comparisons in terms of coupled transcriptomic distance
Static Pair CTF-H OT DeST-OT TOAST MOSCOT
$(1,2)$ $33.58$ $32.64$ $34.13$ $33.07$
$(2,3)$ $18.01$ $14.22$ $21.87$ $15.63$
$(3,4)$ $43.26$ $42.26$ $43.89$ $43.19$
$(4,5)$ $20.14$ $18.47$ $20.54$ $18.40$
@@
Table 9: Comparisons in terms of runtime (s) with varying batch size.
Batch Size CTF-H OT DeST-OT MOSCOT TOAST
$10$ $0.0124$ $0.0114$ $0.5199$ $0.0227$
$50$ $0.0017$ $0.0351$ $0.8479$ $0.0114$
$100$ $0.0011$ $0.1016$ $2.3381$ $0.0167$
$150$ $0.0018$ $0.1778$ $1.1979$ $0.0247$
$200$ $0.0034$ $0.3521$ $1.5328$ $0.0337$
$250$ $0.0025$ $0.3836$ $1.9100$ $0.0426$
$300$ $0.0035$ $0.5323$ $1.0772$ $0.0543$
$350$ $0.0044$ $0.5935$ $1.2983$ $0.0656$
$400$ $0.0047$ $0.6217$ $1.1782$ $0.0817$
$450$ $0.0051$ $0.7806$ $1.2357$ $0.0983$
$500$ $0.0048$ $0.9984$ $1.2243$ $0.1197$
$550$ $0.0054$ $1.3473$ $1.4695$ $0.1468$
$600$ $0.0073$ $1.6192$ $1.6680$ $0.1905$
$650$ $0.0070$ $1.7771$ $2.2559$ $0.2077$
$700$ $0.0090$ $2.2104$ $1.9901$ $0.2309$
$750$ $0.0089$ $2.9324$ $2.7181$ $0.2574$
$800$ $0.0106$ $3.2879$ $1.9840$ $0.3001$
$850$ $0.0117$ $5.0941$ $2.3511$ $0.3229$
$900$ $0.0128$ $5.6585$ $2.4864$ $0.3798$
$950$ $0.0150$ $5.3771$ $3.0964$ $0.4206$
$1000$ $0.0146$ $5.9733$ $3.3683$ $0.4931$
Runtime. To complement Figure 3(d), we report the runtime complexity of the above-mentioned OT methods in Table 9. CTF-H OT is the fastest of the three, followed by TOAST, MOSCOT, and DeST-OT, and is competitive in the metrics above. We also observe that DeST-OT is the slowest, as expected, since its OT objective includes a Gromov-Wasserstein term, which has $O(n^{3})$ runtime, along with other growth- and tissue-distortion-specific terms. To summarize, our evaluations above demonstrate that ContextFlow’s design choices enable it to be highly scalable compared to state-of-the-art methods, while remaining competitive across several spatiotemporal OT alignment metrics.
Figure 5: Visualization of biologically implausible cell type couplings resulted from the PAER-OT plan used in ContextFlow on the Brain Regeneration dataset, where implausibility refers to transitions involving excitatory-inhibitory lineage switches. This result is in sharp contrast to Figure 1 with 55 percent decrease in implausible couplings in this case.
Appendix G Visualizations
@@
To demonstrate the need for integrating biological priors within a generative framework, we compute the EOT plan (Equation 2) for the MOTFM framework and the PAER-OT plan (Section 4.2) for the ContextFlow framework. From these transport plans, we sample couplings corresponding to the first two stages of the Brain Regeneration dataset [Wei et al., 2022] with their associated cell types. Figures 1 and 5 illustrate the excitatory–inhibitory lineage switches present in these couplings. Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
In our transport plan couplings, we observed the following cell type lineage switches: (i) Immature MSN $\rightarrow$ Immature nptxEX, (ii) Immature MSN $\rightarrow$ Immature dpEX, (iii) Immature MSN $\rightarrow$ Immature CMPN, (iv) Immature nptxEX $\rightarrow$ Immature cckIN, and (v) Immature nptxEX $\rightarrow$ Immature MSN. Of these, $54$ implausible transitions arise from the Entropic-OT plan, compared to $24$ from PAER-OT, with the specific transitions detailed in the figure legends. We also observed that the EOT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right. In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information. These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
G.2 Cell Type Distributions across Datasets
@@ G.2.3 Liver Regeneration
G.3 Ligand-Receptor Interactions
Figure 9 illustrates the ligand-receptor score of the NPTX2-NPTXR pair in two consecutive slides from the Brain regeneration dataset. Similar activities are bilaterally visible in the cerebral cortex, suggesting that ligand–receptor interactions are preserved over time and spatially aligned with underlying tissue structure. This observation provides strong evidence that including LR interactions as contextual priors is biologically meaningful, as they capture functional communication signals between cells that remain stable across short time intervals.
Based on the activation of NPTX2–NPTXR in Figure 9, we observe that the communication pattern naturally biases the optimal couplings towards transitions, such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 10). These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context. This example highlights the richness of the contextual information captured by our proposed biological prior and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process yields more interpretable, biologically consistent trajectories.